I am kind of unsure of my response for this equation:
It asks me to rewrite the equation in terms of "w"
A=2(L+W)
This is my work...
A=2(L+W)
A=2L+2W
A-2L=2W
((A-2L)/2)=W --------> I'm not sure if this would be my final answer or could it be (((A/2)-L))=W
Please explain to me why my first solution is right/wrong. Thank you.

Found 2 solutions by solver91311, DrBeeee:
Answer by solver91311(24713)



In the absence of any other instructions your professor/instructor might have given as to the form of the answer, either of your answers should be acceptable. In practical terms, either could end up being preferrable assuming that this process was an intermediate step in a larger problem and depending on what the next step in the problem might be. Be that as it may, your work is spot on and both of the answers you provided are mathematically equivalent and correct. Good job.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by DrBeeee(684)

Either answer is correct. I usually hedge my answer by giving both as follows;
(1) W = (A - 2L)/2 or
(2) W = A/2 - L
and let the reader choose the preferred form.
Note a couple of differences that I have from your answers. The first is that you always put the variable that you want, W in this case, on the LEFT side of the equation. Then you read it as W = something, not something = W. The second is that I omit all of the unneccessary parentheses in the answer. This gives a neater appearance to the answer and easier for the reader to "see" the answer. I also have a comment on the given formula. It is the formula for the perimeter of a rectangle, so I'd expect the standard practice of using the variable p, instead of A, which is usually used for area as in
(3) A = lw.
In addition, lower case letters are used for length and width as in
(4) p = 2l + 2w.
(My preference) Answer: w = p/2 - l